f={(1,2),(2,3),(3,3)},g={(1,0)(2,3),(3,2),(4,8)}
Find the domain of f/g=

Answer by greenestamps(13198)



The possible values in the domain for f/g are the values in the domains of both f and g: {1,2,3}.

4 can't be in the domain of f/g because it is in the domain of g but not of f.

Then observe that (f/g)(1) = 2/0 is undefined; so x=1 is not in the domain of f/g.

ANSWER: The domain of f/g is {2,3}.